OpenStudy (vshiroky):

The value of 3(x+2)^2 when x =-4 I can't remember the order

OpenStudy (anonymous):

remember, PEMDAS P Parentheses first E Exponents (ie Powers and Square Roots, etc.) MD Multiplication and Division (left-to-right) AS Addition and Subtraction (left-to-right)

OpenStudy (vshiroky):

Do I foil it?

OpenStudy (anonymous):

Well you already have what x is, so you would just plug x in. If you were expanding, then yes, you foil it.

OpenStudy (vshiroky):

I got 36 but I think it's wrong

OpenStudy (anonymous):

so since x = -4 \[3(-4+2)^2\]

OpenStudy (anonymous):

sorry, \[3(-2)^2\]

OpenStudy (anonymous):

\[3(-2)^2 = 3(4)\]

OpenStudy (anonymous):

\[3\times4\]

OpenStudy (anonymous):

and that equals 12

OpenStudy (vshiroky):

Thank you can you double check another answer for me

OpenStudy (vshiroky):

9y^2-4x X=-4 and y=1

OpenStudy (usukidoll):

\[9y^2-4x\]

OpenStudy (usukidoll):

and we have y =1, and x = -4 right?

OpenStudy (vshiroky):

I got 25

OpenStudy (usukidoll):

so we plug in 1 and -4 into the equation \[9(1)^2-4(-4)\]

OpenStudy (usukidoll):

\[9(1)-(-16) \rightarrow 9+16=25\]

OpenStudy (usukidoll):

I also got 25, so it seems to be the right answer.

OpenStudy (vshiroky):

Ty!